A river flows due south with a speed of 2.00 {\rm m}/{\rm s} . A man steers a motorboat across the river; hisvelocity relative to the water is 4.20 {\rm m}/{\rm s} due east. The river is 800 {\rm m} wide.
a)   What is the magnitude of his velocityrelative to the earth?
         v=     m/s
b)   What is the direction of his velocityrelative to the earth?
         θ=       o south ofeast
c)   How much time is required for the man tocross the river?
        t=      s
d)   How far south of his starting point willhe reach the opposite bank?
         d=     m
Answer #1
Concepts and reason

The main concept used is the magnitude and the direction of the vector.

Initially, draw the diagram for the system and form the diagram specify xandyx{\rm{ and }}y component of the system. Finally, this component is responsible for the calculation of the magnitude and the direction of the system.

Fundamentals

The expression of the magnitude for xandyx{\rm{ and }}y component is equal to,

r=x2+y2r = \sqrt {{x^2} + {y^2}}
‎Here, rr is the magnitude.

The expression of the direction of the xandyx{\rm{ and }}y is equal to,

θ=tan1(yx)\theta = {\tan ^{ - 1}}\left( {\frac{y}{x}} \right)

Here, θ\theta is the angle between the component xandyx{\rm{ and }}y .

(a)

The diagram for the provided system is given below,

4.2 m/s
2.0 m/s
800 m

From the above diagram, it is clear the velocity in xx and yy component is equal to,

Velocity in yy direction is equal to, vy=2.0m/s{v_y} = 2.0{\rm{ m/s}}

Velocity in xx direction is equal to, vx=4.2m/s{v_x} = 4.2{\rm{ m/s}}

The expression of the magnitude of his velocity relative to the earth is equal to,

v=vx2+vy2v = \sqrt {v_x^2 + v_y^2}

Here, vx{v_x} is the velocity in xx component, vy{v_y} is the velocity in yy component and vv is the resultant velocity.

Substitute 2.0m/s22.0{\rm{ m/}}{{\rm{s}}^2} for vy{v_y} and 4.2m/s24.2{\rm{ m/}}{{\rm{s}}^2} for vx{v_x} in the above expression of the resultant velocity,

v=(4.2m/s2)2+(2.0m/s2)2=4.65m/s\begin{array}{c}\\v = \sqrt {{{\left( {4.2{\rm{ m/}}{{\rm{s}}^2}} \right)}^2} + {{\left( {2.0{\rm{ m/}}{{\rm{s}}^2}} \right)}^2}} \\\\ = 4.65{\rm{ m/s}}\\\end{array}

(b)

Calculation of direction of his velocity relative to the earth.

The expression of the direction in terms of the velocity is equal to,

θ=tan1(vyvx)\theta = {\tan ^{ - 1}}\left( {\frac{{{v_y}}}{{{v_x}}}} \right)

Substitute 2.0m/s22.0{\rm{ m/}}{{\rm{s}}^2} for vy{v_y} and 4.2m/s24.2{\rm{ m/}}{{\rm{s}}^2} for vx{v_x} in the above expression of the resultant velocity,

θ=tan1(2.0m/s24.2m/s2)=25.46\begin{array}{c}\\\theta = {\tan ^{ - 1}}\left( {\frac{{2.0{\rm{ m/}}{{\rm{s}}^2}}}{{4.2{\rm{ m/}}{{\rm{s}}^2}}}} \right)\\\\ = 25.46^\circ \\\end{array}

(c)

Calculate the time required for the man to cross the river.

The expression of the distance in terms of the speed is equal to,

s=vts = vt

Rearrange the above expression in terms of the time tt .

t=svt = \frac{s}{v}

Substitute 800m800{\rm{ m}} for ss and 4.2m/s4.2\;{\rm{m/s}} for vv in the above expression of the time,

t=800m4.2m/s=190.47s\begin{array}{c}\\t = \frac{{800{\rm{ m}}}}{{4.2\;{\rm{m/s}}}}\\\\ = 190.47{\rm{ s}}\\\end{array}

(d)

Calculate the starting point while reach the opposite bank of the river.

The expression of the distance while reach the opposite bank of the river is equal to,

s=vts = vt

Substitute 2.0m/s2.0{\rm{ m/s}} for vv and 190.47s190.47{\rm{ s}} for tt in the above expression of the distance ss

s=(2.0m/s)(190.47s)=380.94m\begin{array}{c}\\s = \left( {2.0{\rm{ m/s}}} \right)\left( {190.47{\rm{ s}}} \right)\\\\ = 380.94{\rm{ m}}\\\end{array}

Ans: Part a

The magnitude of his velocity relative to the earth is equal to 4.65m/s4.65{\rm{ m/s}} .

Part b

The direction of his velocity relative to the earth is equal to 25.4625.46^\circ .

Part c

The time required for the man to cross the river is equal to 190.47s190.47{\rm{ s}} .

Part d

The starting point while reach the opposite bank of the river is equal to 380.94m380.94{\rm{ m}} .